What is equivalent resistance when we connected two resistances in parallel.
1
Expert's answer
2020-10-05T16:11:02-0400

Solution

If there are two resistances "R_1, and\space R_2" in parallel combination in any circuit then equivalent resistance R can be expressed as

"\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}"

Or

"R=\frac{R_1 R_2}{R_1+R_2}"